Grade 11Mechanics

height=171 ii) State the approximation that has to be made in this equation to reduce it to an equation of simpleharmonic motion.

abhishek shekhar
7 Years agoGrade 11

1 Answer

Samyak Jain
ApprovedApproved Tutor Answer7 Years ago
Displace the pendulum slightly from mean (vertical) position making angle \theta with vertical and resolve mg along the string and perpendicular to it.
You will find that the restoring force of simple pendulum is mgsin\theta.
\therefore restoring force is F = – mgsin\theta. Also, F = ma.
a = – gsin\theta.   \because a = d2x / dt2 ,   d2x / dt2 = – gsin\theta        ….(1)
Now, x is slight displacement, considering the circular part of moion, we get
x = l\theta   \Rightarrow  \theta = x / l
Differentiate both sides wrt t i.e. d\theta/dt = (1/l)dx/dt.  Here 1/l is constant, l being length of string.
Again differentiate both sides wrt t to get d2\theta/dt2 = (1/l)d2x/dt2  or  d2x/dt2 = ld2\theta/dt2 .
Put in (1).
ld2\theta/dt2 = – gsin\theta, which is the required equation to be obtained.
sin\theta = \theta is the approximation to be made to reduce above equation to an equation of simple harmonic motion.